Requirement 3 please
Assume Nortel Networks contracted to provide a customer with Internet infrastructure for $2,700,000. The project began in 2018 and was completed in 2019. Data relating to the contract are summarized below: 2018 356,000 $2,135,000 2019 Costs incurred during the year Estinated costs to complete as of 12/31 illings during the year Cash collections during the year 1,424,000 475, 000 1,760,000 278,000 1,820,000 Required: 1. Compute the amount of revenue and gross profit or loss to be recognized in 2018 and 2019 assuming Nortel recognizes revenue over time according to percentage of completion. 2. Compute the amount of revenue and gross profit or loss to be recognized in 2018 and 2019 assuming this project does not qualify for revenue recognition over time. 3. Prepare a partial balance sheet to show how the information related to this contract would be presented at the end of 2018 assuming Nortel recognizes revenue over time according to percentage of completion. 4. Prepare a partial balance sheet to show how the information related to this contract would be presented at the end of 2018 assuming this project does not qualify for revenue recognition over time.
Answer #1
1 2018 2019
Contract price (1) 2700000 2700000
Actual cost to date (a) 356000 2491000
(356000+2135000)
Estimated cost to complete (b) 1424000 0
Total estimated cost ©=(a)+(b) 1780000 2491000
Percentage of completion (a)/© 20% 100%
Expected gross profit (1)-© 920000 209000
Revenue to be recognized=(Contract price*Percentage of completion)-Revenue recognized earlier
Gross profit to be recognized=(Expected gross profit*Percentage of completion)-Gross profit already recognized
2018
Revenue to be recognized=(2700000*20%)-0=$ 540000
Gross profit to be recognized=(920000*20%)-0=$ 184000
2019
Revenue to be recognized=(2700000*100%)-540000=$ 2160000
Gross profit to be recognized=(209000*100%)-184000=$ 25000
2 If the project does not qualify for revenue recognition over time,recognize revenue on completion of contract
2018
Revenue to be recognized=0
Gross profit to be recognized=0
(Since contract has not completed)
2019
Revenue to be recognized=Contract price=$ 2700000
Gross profit to be recognized=Contract price-Actual cost to date=2700000-2491000=$ 209000
3 Partial balance sheet-2018
Current assets:
Accounts receivable
Billings made during the year 475000
Less: Cash collections 278000 197000
Cost and profit in excess of billings
Cost incurred 356000
Add:Gross profit recognized 184000
540000
Less: Billings 475000 65000
4 Partial balance sheet-2018
Current assets:
Accounts receivable
Billings made during the year 475000
Less: Cash collections 278000 197000
Current liability:
Billings in excess of cost
Billings made during the year 475000
Less: Cost incurred 356000 119000